Let (x1, x2) be some interior bundle on the budget constraint, and suppose p1 = 4...

  1. Let (x1, x2) be some interior bundle on the budget constraint, and suppose p1 = 4 and p2 = 2. If the marginal rate of substitution at (x1, x2) is -3, then is (x1, x2) optimal? Explain why or why not, and justify your answer with economic intuition.
Answer #1

At the optimal point,

MRS(x1,x2) = -P1 / P2

Substitute given values,

MRS = -3

-P1/P2 = -4/2 = -2

As it can be seen, MRS is different from the price ratio. This implies the current point is not optimal and the consumer needs to change his consumption of x1 and x2 to reach a point where his MRS equals price ratio